OpenStudy (anonymous):

Samarium has a half-life of approximately 106 billion years. If you found a sample of samarium from the beginning of Earth's history, approximately 4.6 billion years ago, which would be the best approximation of the amount of parent isotope left in the rock? Nearly 100 percent 50 percent 25 percent Less than 10 percent

OpenStudy (aaronq):

106 = ln2/k k= 6.539x10^-3 A=Aoe^(-kt) A=100e^(-6.539x10^-3(4.6)) A= 97.036 %

OpenStudy (anonymous):

1st option.

thomaster (thomaster):

So.. after 106 billion years the sample contains 50% of isotopes 4.6 billion is only like 5% of that so its very logical that the answer is nearly 100%

OpenStudy (goformit100):

Nearly 100 percent is the correct option